Following a rollback of the whole system undertaken by Updata last night, we believe that all connectivity services are now working as expected. If this is not the case for your school, please let us know via the ICT Service Desk on 01438 844777 Option 1 and then 2
Thank you for your patience over the past few days. We are fully aware of how disruptive this must have been to you all, and we will be expecting a full incident report from Updata in due course
Updata have completed the roll back of the service now. However, SSL inspection is currently on for all sites, the result of this is visiting Google or YouTube you will receive a security error. We have asked Updata to revert this as soon as possible. A workaround is to deploy the Updata SSL certificate to your network or standalone device a guide can be found at “ssl-filtering.updata.net”.
We apologise for the continued disruption, a full report will be issued once the incident has been resolved.
Head of Technology Services Development
Good morning all
Unfortunately, there are continuing issues with the HICs network this morning. Most of the issues are fixed but the SSL Inspection problems are still present.
We can only apologise for the continued disruption, and assure you that we are doing everything we can to get the service restored
We believe the transparent proxy is now working as it should be but please contact us if you have any concerns.
It also appears that the schools who hadn’t previously set up SSL inspection had this deployed over the weekend in error. This means that devices which do not have the Updata CA certificate installed would get error messages when accessing certain HTTPS websites. Updata are looking to roll this back but it is proving to be problematic. Updata inform us that this should be done by 8pm this evening and with this in mind we are hopeful that by tomorrow morning, there will be a stable service once again.
Further updates will follow and thanks as ever for your understanding.
Updata after consultation with HICS and HCC, have decided the next step is to revert the changes made to the Netsweeper filtering platform and put everything back as it was prior to the attempted upgrade. They are expecting this work to take 4-5 hours to complete from 13:00 on 26/06/17. HICS/HCC are in close communication with Updata and if we receive any updates between now and the close of business, we will communicate them.
- Transparent proxy network is completely down at all sites within HICS
- SSL Inspection has been turned on at all sites regardless of if it was on previously
- Access to the Netsweeper proxy admin console is unavailable
- WSUS synchronisations are failing
- GSuite/Chromebook connection issues relating to SSL inspection
There may be some localised issues however once the change has been reverted these should be resolved.
We can understand the frustrations this loss of service is causing, rest assured this is being raised to the highest level within HfL/Updata & HCC.